# Memory, Consciousness and Large Language Model

### v1

@@ 1   Introduction
Instead of addressing this question directly, let’s consider it from another perspective: What do we, as conscious human beings, have in common with LLMs? The answer that immediately comes to mind is memory. The study of memories in LLMs is crucial for solving issues like catastrophic forgetting and hallucination. Moreover, there exists a profound connection between memory and consciousness, as elucidated by Tulving’s research in the last century (Tulving, 1985). During our research on memories in LLM and in Tulving’s theory of memory, we have uncovered a series of intriguing coincidences between these two seemingly disparate subjects. Based on these coincidences, we propose a bold conjecture suggesting the presence of a duality between Tulving’s memory theory and various memories in LLM. This duality leads us to attribute the relationship between memory and consciousness in Tulving’s theory as an emergent ability.
The structure of this paper is outlined as follows: In Section 2, we introduce Tulving’s theory of memory and consciousness. In Section 3, we propose a conjecture suggesting the existence of a duality between memories in LLM and those in Tulving’s theory. We provide a detailed explanation of how each type of memory corresponds between these two distinct areas. In Section 4, building upon this duality, we establish a potential correspondence between the synergistic ecphory model of retrieval and emergent ability. Additionally, we present supporting evidence for our argument. In Section 5, we further speculate that consciousness is a form of emergent ability, drawing upon our previous arguments and other corroborating evidence. Finally, in Section 6, we further discuss our consciousness theory.
@@ 3   Duality between LLM and Tulving’s Theory of Memory
Identifying the corresponding memory for episodic memory in LLM is of paramount importance. For episodic memory, an essential aspect has been discussed in Tulving’s work. Tulving(Tulving, 1985; Budson et al., 2022) defined episodic memory as a set of processes that enable us to mentally time-travel and re-experience past moments. These processes involve the initial intake of information from our sensory stores and working memory, followed by the creation of a mental representation of a specific moment in time. In the realm of Natural Language Processing (NLP), we interpret such a description to refer to ”time-series” information. We can no longer retain episodic memories in the same manner as we can with the other two types of memory. This is primarily because the current architecture of LLM cannot store time-series data within its parameters. In other words, when you store a date by fine-tuning, it serves merely as a numerical value for the LLMs, rather than representing a genuine recollection of the past.
@@
Similar attempts to establish a correspondence between memory in psychology and memory in LLM have been made in many articles about LLM agents (Weng, 2023). However, their arguments are based on intuition and the assumption of the current limitation in LLM memory. It is another ”bitter lesson”(Sutton, 2019) happening right now.
@@ 4   Synergistic-Ecphory-Model/Emergent-Ability Correspondence
The blue line in Fig.(0(a)) has the same horizontal axis value, indicating that we are using identical trace information. In the context of LLM, this implies the application of a specific prompting strategy. Then we increase the vertical axis value from A1 to D1, signifying an increase in either the model’s size or the amount of training data. As we transition from A1 to B1, we remain within Region 1, where the ecphoric information (emergent ability) remains unattained. Progressing from B1 to C1, we enter Region 2 and attain some familiarity with ecphoric information, although it remains insufficient to fully realize this ability. Consequently, we observe an improvement in performance as the model size increases. Transitioning from C1 to D1 places us in Region 3, where the ecphoric information (emergent abilities) is fully achieved, rendering further improvement unlikely. We draw the performance curve on the right panel of Fig.(0(a)), which aligns with performance curves of most emergent abilities. Similarly, we observe a correspondence between the SEM and in-context learning in LLM, as illustrated in Fig.(0(b)). This correspondence also explains why LLM can function as a few-shot learner, as demonstrated in the plot. The same deduction can be applied to other few-shot prompting methods, such as the chain of thought.
Here we provide some evidence that supports our conjecture regarding this correspondence. It has been demonstrated that emergent performance does not manifest in the absence of in-context learning (Lu et al., 2023). This distinction arises because the ecphoric process requires both trace information (input context) and retrieval information (model size/quality). Additionally, it has been observed that Large Language Models (LLMs) do not learn new tasks during test time (Min et al., 2022). Their analysis shows that the model may ignore the task defined by the demonstrations and instead use prior from pertaining. This phenomenon can be better understood in the view of the ecphory process. Furthermore, we have noticed that several other papers align with our conjecture: (Wang et al., 2023; Chan et al., 2022).
@@ 5   Consciousness as an Emergent Ability
So far, all the emergent abilities we have observed have originated from LLMs with substantial model sizes and extensive training data. However, if our theory regarding the correlation between the SEM of retrieval and the emergent abilities of LLMs is accurate, we should also anticipate witnessing emergent abilities generated by smaller models with significantly extended context lengths, as illustrated in Fig.(2). We firmly believe that such emergent abilities do indeed exist, and one plausible candidate among them is consciousness. (In this section, consciousness refers to the widely discussed general consciousness, closely related to Tulving’s concept of autonoetic consciousness. )

### v2

@@ 1   Introduction
Instead of addressing this question directly, let’s consider it from another perspective: What do we, as conscious human beings, have in common with LLMs? The answer that immediately comes to mind is memory. The study of memories in LLMs is crucial for solving issues like catastrophic forgetting and hallucination. Moreover, there exists a profound connection between memory and consciousness, as elucidated by Tulving’s research in the last century (Tulving, 1985). During our research on memories in LLM and in Tulving’s theory of memory, we have uncovered a series of intriguing coincidences between these two seemingly disparate subjects. Thus we propose a bold conjecture suggesting the presence of a duality between Tulving’s memory theory and various memories in LLM. Expanding upon this duality, we further identify a potential correspondence between Tulving’s synergistic ecphory model (SEM) of retrieval and the emergent abilities observed in LLMs. This perspective offers a novel approach to comprehending emergent abilities and in-context learning. This theory is consistent with several existing experimental observations, including (Lu et al., 2023; Min et al., 2022; Wang et al., 2023; Chan et al., 2022; Wei et al., 2022; Brown et al., 2020). The duality finally leads us to attribute the relationship between memory and consciousness in Tulving’s theory as an emergent ability.
The structure of this paper is outlined as follows: In Section 2, we introduce Tulving’s theory of memory and consciousness. In Section 3, we propose a conjecture suggesting the existence of a duality between memories in LLM and those in Tulving’s theory. We provide a detailed explanation of how each type of memory corresponds between these two distinct areas. In Section 4, building upon this duality, we establish a potential correspondence between the synergistic ecphory model of retrieval and emergent ability. Additionally, we present supporting evidence for our argument. In Section 5, we speculate that consciousness is a form of emergent ability, drawing upon our previous arguments and other corroborating evidence. In Section 6, we further discuss our consciousness theory in detail. Finally, the conclusion and further study are given in Section 7.
@@ 3   Duality between LLM and Tulving’s Theory of Memory
Identifying the corresponding memory for episodic memory in LLM is of paramount importance. For episodic memory, an essential aspect has been discussed in Tulving’s work. Tulving(Tulving, 1985; Budson et al., 2022) defined episodic memory as the set of processes that enable us to mentally time-travel and re-experience past moments. These processes involve the initial intake of information from our sensory stores and working memory, followed by the creation of a mental representation of a specific moment in time. In the realm of Natural Language Processing (NLP), we interpret such a description to refer to ”time-series” information. We can no longer retain episodic memories in the same manner as we can with the other two types of memory. This is primarily because the current architecture of LLM cannot store time-series data within its parameters. In other words, when you store a date by fine-tuning, it serves merely as a numerical value for the LLMs, rather than representing a genuine recollection of the past.
@@
Similar attempts to establish a correspondence between memory in psychology and memory in LLM have been made in many articles about LLM agents (Weng, 2023). However, their arguments are based on the current limitation in context length, and therefore they require an external vector database for long-term memory. It is another ”bitter lesson”(Sutton, 2019) happening right now.
@@
Table 2: A schematic diagram of the relation between memory systems from Tulving’s theory and different memories in LLMs.
@@ 4   Synergistic-Ecphory-Model/Emergent-Ability Correspondence
The blue line in Fig.(1(a)) has the same horizontal axis value, indicating that we are using identical trace information. In the context of LLM, this implies the application of a specific prompting strategy. Then we increase the vertical axis value from A1 to D1, signifying an increase in either the model’s size or the amount of training data. As we transition from A1 to B1, we remain within Region 1, where the ecphoric information (emergent ability) remains unattained. Progressing from B1 to C1, we enter Region 2 and attain some familiarity with ecphoric information, although it remains insufficient to fully realize this ability. Consequently, we observe an improvement in performance as the model size increases. Transitioning from C1 to D1 places us in Region 3, where the ecphoric information (emergent abilities) is fully achieved, rendering further improvement unlikely. We draw the performance curve on the right panel of Fig.(1(a)), which aligns with performance curves of most emergent abilities. Similarly, we observe a correspondence between the SEM and in-context learning in LLM, as illustrated in Fig.(1(b)). A large model contains a large amount of semantic information, which means only a little episodic information is needed to reach the familiarity threshold. It explains why LLM can function as a few-shot learner. The same deduction can be applied to other few-shot prompting methods, such as the chain of thought.
Here we provide some existing experimental observations that match our theory regarding this correspondence.
  • (Wei et al., 2022; Brown et al., 2020) The SEM successfully explained the shape of the performance curve for emergent ability and in-context learning. Most importantly, it accounted for both factors simultaneously.
  • (Lu et al., 2023) It has been demonstrated that emergent performance does not manifest in the absence of in-context learning in this work. This phenomenon can be perfectly explained by the SEM that the ecphoric process requires both trace information (input context) and retrieval information (model size/quality).
  • Min et al. (2022) It has been observed that LLMs do not learn new tasks during test time. Their analysis shows that the model may ignore the task defined by the demonstrations and instead use prior from pertaining. This phenomenon can be better understood in the view of the ecphory process that the abilities are actually ”recalled” from the model.
Furthermore, we have noticed that several other papers align with our conjecture: (Wang et al., 2023; Chan et al., 2022).
@@ 5   Consciousness as an Emergent Ability
So far, all the emergent abilities we have observed have originated from LLMs with substantial model sizes and extensive training data. However, if our theory regarding the correlation between the SEM of retrieval and the emergent abilities of LLMs is accurate, we should also anticipate witnessing emergent abilities generated by relatively smaller models with significantly extended context lengths, as illustrated in Fig.(2). We firmly believe that such emergent abilities do indeed exist, and one plausible candidate among them is consciousness. (In this section, consciousness refers to the widely discussed general consciousness, closely related to Tulving’s concept of autonoetic consciousness.)